Write a line in Slope-intercept form that is parallel to 5y-60+8x and contains
(-6,-10)

Answer by mananth(16946)

5y-60+8x
what is the equation
is it 5y-60+8x=0
is it 5y-60=8x
OR
is it 5y=60+8x